Two consecutive odd numbers are that seven times the smaller , subtracted from nine times the bigger gives 144. Find the two numbers

Answer by ikleyn(53765)

.

Let the numbers be  n and  (n+2)  with an odd integer n.


An equation is


    9*(n+2) - 7n = 144.


Simplify and solve


    9n + 18 - 7n = 144

    2n           = 144 - 18

    2n           = 126

     n           = 126/2 = 63.   


ANSWER.  The numbers are 63 and 65.

Solved.